If f(x) = x^2 + 1 and g(x) = x - 2, find[f degrees g ](x)
a. x^2 - 4x + 5
b. x^2 - 1
c. x^2 - 3
d. x^3 - 2x^2 + x - 2

Answer by user_dude2008(1861)

(x-2)^2+1 = x^2-4x+4+1 = x^2-4x+5
Answer: a) x^2 - 4x + 5